OpenStudy (ondinana):

Help please...! What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4? y =____x +_____

OpenStudy (ondinana):

@paki @phi @Compassionate ??? Some please help, I'll medal. (And fan unless I already have.)

OpenStudy (amorfide):

the equation of a line \[y-y _{1}=m(x-x _{1})\] where your set of coordinates are \[(-1,5)=(x _{1},y _{1})\]

OpenStudy (amorfide):

m is the slope of the line

OpenStudy (texaschic101):

couldn't you just use y = mx + b ? subbing in -4 for m (slope) and your points and just solve for b, the y intercept ?

OpenStudy (amorfide):

yeah you could

OpenStudy (amorfide):

much simpler lol

OpenStudy (ondinana):

So would it be..; 5=___-1=___? I'm so confused I don't understand.

OpenStudy (amorfide):

you have y=mx+c for the standard equation of a line you know y=5 x=-1 and m=-4 sub it in 5=-4(-1)+c 5=4+c work out c

OpenStudy (amorfide):

credit goes to @texaschic101

OpenStudy (texaschic101):

you could have used either formula and you would get the same answer

OpenStudy (ondinana):

Okay so.. 5=-4+c and I'll solve it and can you check it for me?

OpenStudy (ondinana):

Blah typo oh well

OpenStudy (texaschic101):

what did you get ?

OpenStudy (ondinana):

I got 1?

OpenStudy (amorfide):

well done

OpenStudy (texaschic101):

so your equation is : y = -4x + 1....because that number you just solved for was the y intercept

OpenStudy (texaschic101):

and you already knew the slope (m) was -4

OpenStudy (ondinana):

Oh okay! I sorta get it all now, thank you both!! :D

OpenStudy (texaschic101):

:)